OpenStudy (anonymous):

If f(x)=√(x+3)) how do i set up f(x+Δx)

OpenStudy (anonymous):

\[f \left( x+\Delta x \right)=\sqrt{\left( x+\Delta x+3 \right)}\]

OpenStudy (anonymous):

i have no idea what i do with the delta sign

OpenStudy (anonymous):

Let me know what you want to do.

OpenStudy (anonymous):

simplify I've just never had a delta sign so it has me confused on what i can do with it

OpenStudy (anonymous):

\[\Delta x is a \small increment \in x\] write x+delta x in place of x

OpenStudy (anonymous):

so is √((x+Δx)+3) the simplest it will get?